# Word and Phrase Features in Graph Convolutional Network for Automatic Question Classification

### v2

@@ Abstract
Effective question classification is crucial for AI-driven educational tools, enabling adaptive learning systems to categorize questions by skill area, difficulty level, and competence. This classification not only supports educational diagnostics and analytics but also enhances complex tasks like information retrieval and question answering by associating questions with relevant categories. Traditional methods, often based on word embeddings and conventional classifiers, struggle to capture the nuanced relationships in natural language, leading to suboptimal performance. To address this, we propose a novel approach leveraging graph convolutional networks, named Phrase Question-Graph Convolutional Network (PQ-GCN) to better model the inherent structure of questions. By representing questions as graphs—where nodes signify words or phrases and edges denote syntactic or semantic relationships—–our method allows the model to learn from the interconnected nature of language more effectively. Additionally, we explore the incorporation of phrase-based features to enhance classification performance on question datasets of various domains and characteristics. Our findings demonstrate that the proposed model, augmented with these features, offer a promising solution for more robust and context-aware question classification, bridging the gap between graph neural network research and practical educational applications of AI.
Index Terms:
natural language processing, graph convolutional network, question classification
11footnotetext: Equal contribution
I Introduction
@@
Question classification is vital in providing AI-driven education and training, assisting adaptive learning systems by categorizing questions into skill area or level of difficulty and competence, and providing educational diagnosis and analytics. Accurate automatic question classification also provides headway to more complex tasks such as information retrieval and question answering, in which the association of the given question with a certain category allows for retrieval of relevant contextual knowledge to formulate compatible answers. Traditional methods, relying on word embeddings and conventional classifiers, often struggle with capturing the nuanced relationships between words and phrases, leading to suboptimal classification performance. The complexity of natural language, especially in the context of educational and domain-specific questions, requires more sophisticated approaches that can understand and leverage the inherent structure of language.
@@
Aside from text embeddings, graphs have recently been highlighted as a way to represent unstructured text data. Questions can be naturally represented as graphs, where nodes correspond to words or phrases, and edges capture the relationships between them, such as syntactic dependencies, semantic similarities, or proximity measures. This graph-based representation aligns perfectly with the strengths of graph convolutional network (GCN), which excels at learning from structured data. GCNs are designed to operate on graph-structured data, making them ideal for processing and understanding the interconnected nature of words and phrases within questions. This capability enables GCNs to capture the underlying structure and dependencies in a question, leading to more accurate and context-aware classification decisions. GCNs offer a promising solution by modeling the relationships between words, phrases, and their syntactic and semantic connections as graphs. This approach allows for a more nuanced and context-aware classification, bridging the gap between basic AI research and practical applications in education and beyond.
@@
Question classification poses a unique problem even among the different types of text classification, as the amount of information available in a question tends to be limited compared to document-level text counterparts. While neither tackling the task of text classification with GCNs, nor automated question classification via artificial intelligence is new, our investigation aims to determine the feasibility of adapting GCNs for question classification. Up till our work, GCNs have not been thoroughly explored in the context of question classification, with the exception of [1] and [2] which are reviewed in Section II-A. In particular, we explore additional phrase-based features in our proposed model, PQ-GCN, to enhance classification performance, and assess the effectiveness of these features in low-resource settings.
II Related Work
II-A Automatic Question Classification
@@
Following these works, a review of automatic question classifiers from 2012 to 2017 [6] revealed that there was an increase in the number of works proposing neural network-based methods, or using neural networks in combination with rule-based methods. The approaches include concatenating pre-trained word embeddings to form a question embedding [7], incorporating subword information in character-based convolutional neural network (CNN) [8], and combining CNN and long short-term memory (LSTM) in a hybrid model [9]. These neural networks, effectively balancing feature extraction and sequence modeling, were shown to outperform traditional machine learning techniques and earlier deep learning models in classifying questions across various domains.
While there are other works which focus on question classification using neural networks, we review the two works up to date that specifically makes use of text graphs and highlight the improvements we propose with PQ-GCN. The first work [1] incorporates dependency parsing between words, pointwise mutual information (PMI) between regular expressions, cosine similarity between phrase embeddings, and KL divergence between topics in a heterogeneous graph for question and short document classification. The authors based the architecture on TensorGCN [10] and 4 heterogeneous node sets and adjacency matrices, while PQ-GCN is based on SHINE [11] and uses 5 homogeneous node sets and adjacency matrices which reduces sparsity in input data. The second work [2] uses word graph based on dependency parsing, and trains a graph attention network (GAT) [12]. This approach lacks consideration for other textual features, such as semantic and sequential relationships between words and phrases, that would enable better understanding of a question, given their short lengths.
II-B Text Classification Using Text Graphs
@@
There are also combinatory methods based on GAT. HyperGAT (Hypergraph Attention Networks) [20] builds hypergraphs for each document to capture high-level interaction between words, consisting of sequential hyperedges connecting all words in a sentence and semantic hyperedges connecting top-K words with LDA-based topic modeling. Attention mechanism is utilized at both node- and edge-level to update representations. Contrastive learning has also been utilized in conjunction with topic modeling and GAT to improve classification performance on short texts [21].
III Why Use Phrase Features?
As with all text classification tasks, feature extraction from questions poses a particular challenge, as the features need to be representative of the questions and yet be domain-agnostic to be generalizable across different classification frameworks. We identify a research gap in adapting GCNs for question classification in the lack of question classification-specific feature extraction and graph construction methods. While works discussed in the previous section discuss word and document relationships, phrase relationships have not been explored in the context of text classification with graph representations. The effectiveness of phrases in providing deeper semantic connections in the text has been discussed in topic modeling for questions [22, 1], as well as in graph-based document filtering tasks [23]. This section outlines the rationale to explore phrase relationships as a crucial feature for question classification.
Figure 1: Corpus-level graph construction with extracted features
Capturing Meaningful Chunks of Information
Unlike individual words, phrases (e.g. noun phrases and verb phrases) often encapsulate more specific and contextually relevant information. Consider the question “How do enzymes function in the human body?” The phrase “enzymes function” and “human body” are pivotal in identifying this as a biological process question, likely classified under “Biology” or “Physiology.” The phrases indicates the action or process in question, and also specifies the domain of interest. In the question “Describe the steps to solve a quadratic equation,” the phrases “steps to solve” and “quadratic equation” are linked by their relevance to problem-solving in mathematics. This also applies to question analysis with part-of-speech (POS) tags. POS tags help in identifying the grammatical structure, but they do not capture the full semantic context when observed individually.
Intra-Question Relationships
By extracting phrases and analyzing their semantic embeddings, we can understand the relationship between different components of the question. For instance, in “What are the primary causes of the French Revolution?”, the proximity and relationship between “primary causes” and “French Revolution” are key to categorizing this question under “History” or “Political Science.”
Disambiguation and Contextualization
Phrases provide additional context that can help disambiguate words that might have multiple meanings. For instance, in “What is the role of the cell in an electric circuit?”, the phrase “role of the cell” and its proximity to “electric circuit” helps to clarify that “cell” refers to a component of an electric circuit rather than a biological cell.
@@ III Why Use Phrase Features?
While word-level features such as embeddings and POS tags are foundational for understanding the basic structure and meaning of a question, phrase-level features offer a more comprehensive and contextually rich understanding. By extracting noun phrases and verb phrases and analyzing their semantic embeddings and relationships, one can capture the nuanced meaning and intent of questions, leading to more accurate and effective classification models.
IV Methodology
IV-A Feature Extraction and Graph Construction
@@
To construct meaningful graph representations from a corpus of question text, we select several feature extraction techniques, each contributing uniquely to capturing different aspects of the textual data. A summary of the extracted features are provided in Table I.
TABLE I: Summary of Extracted Features
@@
Figure 2: Overview of model architecture
@@
TABLE II: PQ-GCN Model Parameters
@@
TABLE III: Classification Label Distributions
@@
TABLE IV: CNN Model Parameters
@@
TABLE V: Bi-LSTM Model Parameters
@@
Phrase-Level POS Tags
[…]
Named Entity Recognition (NER)
TABLE VI: Baseline comparisons for macro-averaged F1, Precision, and Recall scores across datasets
Best F1, precision, and recall scores are in bold and the second best scores are underlined for each dataset.
Model Metric NU ARC LREC Bloom TREC
CNN F1 0.085 0.226 0.198 0.752 0.782
@@
TextGCN [13] F1 0.722 0.694 0.671 0.663 0.730
@@
IV-B Model Architecture
As we identify phrase-based features as the gap towards building a more robust automatic question classification model, we also identify a model architecture capable of incorporating phrase-based features as a modular add-on, as well as consolidating each graph of different node and edge types into a single question embedding. Initial approach was to concatenate the outputs of 5 individual 2-layer GCNs, followed by a linear layer and softmax function. Various combinations of 2D convolution layers and pooling layers have also been experimented with, but ultimately we found that the final question embedding was not able to provide sufficient information for classification and performed worse than baseline models.
Following the initial exploration, we base our output layers largely on SHINE [11]. An overview of model architecture is provided in Figure 2. Each graph is passed through their own 2-layer GCN. Each GCN output is then concatenated to form question embeddings, and a question graph is created by using these question embeddings as node features and calculating the cosine similarity between the question embeddings for edge features. Hence, a dynamic graph of question nodes and cosine similarity edges is learned from the individual 2-layer GCN outputs, and this dynamic graph is propagated through a final 2-layer GCN and a linear layer for label classification. We provide the model and training parameters for the proposed PQ-GCN in Table II.
V Experiments
V-A Datasets
We first select a number of datasets used for question classification task to evaluate our proposed model on. The details are outlined in this subsection.
NU
ARC
LREC
Bloom
TREC
Aside from TREC, for which the dataset split is obtained from HuggingFace  [^4] https://huggingface.co/datasets/CogComp/trec, the train/test splits are determined by a fixed seed and can be provided upon request. Relevant statistics for the datasets are found in Table III.
V-B Performance Evaluation
TABLE VII: Keyword Analysis for Five Datasets
Frequency is calculated by % of questions in the label set that contains the keyword.
@@ V-B Performance Evaluation
We compare the performance of our proposed model against selected baseline models in Table VI, including vanilla CNN and Bi-LSTM for non-graph classification models.
The baseline models were selected by reviewing their contributions in graph neural networks for text classification [32], each using distinct feature engineering and graph construction methods for text data, as in Section II, as compared to works which have proposed architectural improvements for computational efficiency. With the exception of CNN and Bi-LSTM, for which parameters were set as outlined in Table IV and Table V, the rest of the baseline models have their model and training parameters set as provided by authors in the project repositories.
PQ-GCN achieves the best macro-averaged F1 scores across most datasets, including NU (0.724), LREC (0.751), and TREC (0.801), while consistently delivering high precision and recall. The improvement from the base SHINE model is also highlighted, as we have kept the original features as proposed in [11], showing the effectiveness of phrase-based features in improving question classification capabilities.
CNN model trained in this experiment is a relatively simple model with no specialized components, yet performs well on Bloom and TREC. This can be attributed to the characteristics of the two corpus, where presence of certain words indicate a strong association to a category. An analysis of most frequently occurring words have been included in Table VII, showing that the Bloom and TREC datasets contains distinct and mutually exclusive class-specific keywords. For example, Bloom’s Synthesis class contains high percentages of command words such as “suggest”, “propose”, and “design” that does not appear in questions of other classes. While the convolution kernels of CNN are effective in capturing these keywords as local features, it performs poorly on NU, ARC, and LREC where there are several overlapping non-class-specific keywords such as “would”, “water”, and “energy”.
@@
TensorGCN performs with the highest F1 score on TREC. TensorGCN obtains its semantic embeddings by training an LSTM model on the fly during graph construction, and this method might be effective in generating more context-accurate embeddings with larger datasets. While this brings about a 0.004 point gain in F1 score over PQ-GCN, it also results in a longer graph construction time with increasing dataset size due to increased training time of the LSTM model, as compared to the proposed methodology of using pre-trained embedding models.
@@
Bi-LSTM has a consistently average performance across the datasets, which is expected of a simple sequence model. TextGCN shows apt results for NU and ARC datasets where its F1 scores are second only to PQ-GCN, but there is room for improvement in leveraging text features other than PMI and TF-IDF when constructing the graph, especially when it comes to text with niche characteristics such as questions. Text-Level GNN shows inconsistent poor performance across the datasets, as its learnable word-word edge weights are highly dependent on relative proximity of keywords and the ability of the message passing mechanism to disambiguate polysemous words based on word proximity. HyperGAT also shows inconsistent performance, possibly attributed to its word-word edge features, which are extracted by segmenting a text at sentence level and treating each sentence as a “window”, as opposed to a sliding window in PMI calculation. Questions in the datasets usually consist of one or two sentences, and lacks usual word-sentence-document hierarchy that HyperGAT leverages on. ME-GCN’s embedding distance-based edge features and InducT-GCN’s text construction that mimics that of TextGCN also do not consider question-specific features, and PQGCN outperforms both models in F1 scores across all datasets.
VI Conclusion

### v3

# PQ-GCN: Enhancing Text Graph Question Classification with Phrase Features

@@ Abstract
Effective question classification is crucial for AI-driven educational tools, enabling adaptive learning systems to categorize questions by skill area, difficulty level, and competence. It not only supports educational diagnostics and analytics but also enhances complex downstream tasks like information retrieval and question answering by associating questions with relevant categories. Traditional methods, often based on word embeddings and conventional classifiers, struggle to capture the nuanced relationships in question statements, leading to suboptimal performance. We propose a novel approach leveraging graph convolutional networks, named Phrase Question-Graph Convolutional Network (PQ-GCN). Through PQ-GCN, we evaluate the incorporation of phrase-based features to enhance classification performance on question datasets of various domains and characteristics. The proposed method, augmented with phrase-based features, outperform baseline graph-based methods in low-resource settings, and performs competitively against language model-based methods with a fraction of their parameter size. Our findings offer a possible solution for more context-aware, parameter-efficient question classification, bridging the gap between graph neural network research and its educational applications.
1 Introduction
@@
Question classification is vital in providing AI-driven education and training, assisting adaptive learning systems by categorizing questions into skill area or level of difficulty and competence, and providing educational diagnosis and analytics. Accurate automatic question classification also provides headway to more complex tasks such as information retrieval and question answering, in which the association of the given question with a certain category allows for retrieval of relevant contextual knowledge to formulate compatible answers.
Traditional methods, relying on word embeddings and conventional classifiers, often struggle with capturing the nuanced relationships between words and phrases, leading to suboptimal classification performance. The complexity of natural language, especially in the context of educational and domain-specific questions, requires more sophisticated approaches that can understand and leverage the inherent structure of language.
@@
Aside from text embeddings, graphs have recently been highlighted as a way to represent unstructured text data. Questions can be naturally represented as graphs, where nodes correspond to different text features, and edges capture the relationships between them, such as syntactic dependencies, semantic similarities, or proximity measures. This graph-based representation aligns perfectly with the strengths of Graph Convolutional Networks (GCNs), which excels at learning from structured data. GCNs are designed to operate on graph-structured data, making them ideal for processing and understanding the interconnected nature of text features within questions. This capability enables GCNs to capture the underlying structure and dependencies in a question, leading to more accurate and context-aware classification decisions. GCNs offer a promising solution by modeling the relationships between text features and their syntactic and semantic connections as graphs, which allows for a more nuanced and context-aware classification.
@@
Question classification poses a unique problem even among the different types of text classification, as the amount of information available in a question tends to be limited compared to document-level text counterparts. While neither tackling the task of text classification with GCNs, nor question classification via automated systems is new, our investigation aims to determine the feasibility of adapting GCNs for question classification. Up till our work, GCNs have not been thoroughly explored in the context of question classification, with the exception of [23] and [20] which are reviewed in Section 2.1. In particular, we explore additional phrase-based features in our proposed model, PQ-GCN, to enhance classification performance, and assess the effectiveness of these features in low-resource settings, where there is limited labeled data available.
2 Related Work
2.1 Automatic Question Classification
@@
Figure 1: Corpus-level graph construction with extracted features. The various features extracted from each question are consolidated at the corpus level to form each feature graph.
Following these works, a review of automatic question classifiers from 2012 to 2017 [22] revealed that there was an increase in the number of works proposing neural network-based methods, or using neural networks in combination with rule-based methods. The approaches include concatenating pre-trained word embeddings to form a question embedding [11], incorporating subword information in character-based Convolutional Neural Network (CNN) [21], and combining CNN and Long Short-Term Memory (LSTM) in a hybrid model [35]. These neural networks, effectively balancing feature extraction and sequence modeling, were shown to outperform traditional machine learning techniques and earlier deep learning models in classifying questions across various domains. More recently, Pre-trained Language Models (PLMs) and Large Language Models (LLMs) which have been trained on large amounts of text data have shown their success in natural language understanding-based tasks such as question generation and classification [2]), but questions might have domain-specific terminologies or technical jargon which may not be captured fully by such generalized models.
While there are other works which focus on question classification using neural networks, we review the two works up to date that specifically makes use of text graphs and highlight the improvements we propose with PQ-GCN. The first work [23] incorporates dependency parsing between words, Pointwise Mutual Information (PMI) between regular expressions, cosine similarity between phrase embeddings, and KL divergence between topics in a heterogeneous graph for question and short document classification. The authors based the architecture on TensorGCN [18] and 4 heterogeneous node sets and adjacency matrices, while PQ-GCN is based on SHINE [31] and uses 5 homogeneous node sets and adjacency matrices which reduces sparsity in input data. The second work [20] uses word graph based on dependency parsing, and trains a Graph Attention Network (GAT) [25]. This approach lacks consideration for other textual features, such as semantic and sequential relationships between words and phrases, that would enable better understanding of questions.
2.2 Text Classification Using Text Graphs
@@
There are also combinatory methods based on GAT. HyperGAT (Hypergraph Attention Networks) [10] builds hypergraphs for each document to capture high-level interaction between words, consisting of sequential hyperedges connecting all words in a sentence and semantic hyperedges connecting top-K words with LDA-based topic modeling. Attention mechanism is utilized at both node- and edge-level to update representations.
While these graph-based methods focus on general document-level text classification, we identify a research gap in adapting them for question classification, with the lack of question-specific feature extraction and graph construction methods, given questions’ comparatively shorter lengths and interrogative sentence structure.
3 Why Use Phrase Features?
As with all text classification tasks, feature extraction from questions poses a particular challenge, as the features need to be representative of the questions’ characteristics and yet be domain-agnostic to be generalizable across different classification frameworks.
There is growing agreement in literature that our mental lexicon contains formulaic language, including idioms, phrases, and multi-word expressions [8], and there are positive benefits from understanding text through parsing by larger, more meaningful chunks rather than individual words [1]. There is sufficient linguistic basis to explore beyond word-level representations in improving natural language understanding capabilities of automatic systems. While works discussed in the previous section discuss word and document relationships, phrase relationships have not been explored in the context of text classification with graph representations. Li and Roth [17] discuss the use of non-overlapping chunks, extracted via a trained classifier, but not explicitly defined phrases.
Parsing by phrases and extracting phrase features have already been proven effective in some areas of computational linguistics, specifically statistical machine translation [16], neural machine translation [30], and topic modeling for questions [24, 23].
@@ 3 Why Use Phrase Features?
While word-level features such as embeddings and POS tags are foundational for understanding the basic structure and meaning of a question, phrase-level features offer a more comprehensive approach to understanding the question text, that can capture meaningful, disambiguated chunks of information. By extracting noun phrases and verb phrases and analyzing their semantic embeddings and relationships, one can capture the nuanced meaning and intent of questions, leading to more accurate and effective classification models.
4 Methodology
4.1 Feature Extraction and Graph Construction
@@
To construct meaningful graph representations from a corpus of question text, we select several feature extraction techniques, each contributing uniquely to capturing different aspects of the textual data. A summary of the extracted features are provided in Table 1.
Table 1: Summary of Extracted Features. Node features for Word POS and Phrase POS are obtained via one-hot encoding, as no viable pre-trained embeddings are available for POS tags.
@@
Figure 2: Overview of model architecture. Each graph is passed through their own 2-layer GCN. Each GCN output is then concatenated to form question embeddings, and a question graph is created by using these question embeddings as node features and calculating the cosine similarity between the question embeddings for edge features. Hence, a dynamic graph of question nodes and cosine similarity edges is learned from the individual 2-layer GCN outputs, and this dynamic graph is propagated through a final 2-layer GCN and a linear layer for label classification.
@@
Table 2: PQ-GCN Model Parameters. Total Parameter Size varies according to number of label classes and total number of unique word-level and phrase-level POS tags (which determines embedding size via one-hot encoding), which differ in each dataset. We report the maximum parameter size of the model that is evaluated on TREC dataset.
@@
Phrase-Level POS Tags
[…]
Named Entity Recognition (NER)
4.2 Model Architecture
As we develop phrase-based features to bridge the gap towards building a more robust automatic question classification model, we also identify a model architecture capable of incorporating phrase-based features as a modular add-on, as well as consolidating each graph of different node and edge types into a single question embedding. Initial approach was to concatenate the outputs of 5 individual 2-layer GCNs, followed by a linear layer and softmax function. Various combinations of 2D convolution layers and pooling layers have also been experimented with, but ultimately we found that the final question embedding was not able to provide sufficient information for classification and performed worse than baseline models.
Following the initial exploration, we base our output layers largely on SHINE [31]. An overview of model architecture is provided in Figure 2, and after a parameter sweep across different layer types, we provide the model and training parameters for the proposed PQ-GCN in Table 2. A training iteration is described in Algorithm 1
Algorithm 1 Pass-through Algorithm for PQ-GCN
1:$\{(V^{(i)},A^{(i)})\}_{i=1}^{5}$: Node sets and adjacency matrices for constructed graphs
2:$\text{GCN}_{i}$: Two-layer GCN for the $i$-th graph
3:$\text{GCN}_{Q}$: Two-layer GCN for question-level graph
4:$f_{\text{linear}}$: Final linear layer for classification
5:Predicted labels for each question
6:for $i=1$ to $5$ do
7:     $H^{(i)}\leftarrow\text{GCN}_{i}(V^{(i)},A^{(i)})$
8:end for
9:$H_{Q}\leftarrow\text{Concat}(H^{(1)},H^{(2)},\dots,H^{(5)})$
10:$A_{Q}\leftarrow\text{CosineSimilarity}(H_{Q},H_{Q})$
11:$Z_{Q}\leftarrow\text{GCN}_{Q}(H_{Q},A_{Q})$
12:$\hat{Y}\leftarrow f_{\text{linear}}(Z_{Q})$
The intermediate states $H^{(i)}$, $H_{Q}$, $A_{Q}$, and $Z_{Q}$ refer to GCN output for graph $i$, question embeddings, question-level dynamic adjacency matrix, and GCN output for question-level graph respectively.
Table 3: Classification label distributions across evaluation datasets
@@ 4.2 Model Architecture
5 Experiments
5.1 Datasets
We first select a number of datasets used for question classification task to evaluate our proposed model on.
NU
ARC
LREC
Bloom
TREC
Aside from TREC, for which the dataset split is obtained from HuggingFace  [^7] https://huggingface.co/datasets/CogComp/trec, the train/test splits are determined by a fixed seed and can be provided upon request. Relevant statistics for the datasets are found in Table 3.
5.2 Baseline Models
We then compare the performance of our proposed model against selected baseline models in Table 6. We include vanilla CNN and Bi-LSTM (with parameters in Table 4 and Table 5) for non-graph classification models; fine-tuned BERT (bert-base-uncased [^8] https://huggingface.co/google-bert/bert-base-uncased) for pre-trained language model. The baseline graph models were selected by reviewing their contributions in graph neural networks for text classification [28], each using distinct feature engineering and graph construction methods for text data, as in Section 2, as compared to works which have proposed architectural improvements for computational efficiency. With the exception of CNN, Bi-LSTM, and BERT, the rest of the baseline models have their model and training parameters set as provided by original authors in their project repositories.
Table 4: CNN Model Parameters
@@ 5.2 Baseline Models
Table 5: Bi-LSTM Model Parameters
@@
Table 6: Baseline comparisons for macro-averaged F1, Precision, and Recall scores across datasets. Best F1 scores are in bold and the second best F1 scores are underlined for each dataset.
Type Model Metric NU ARC LREC Bloom TREC
Non-graph models CNN F1 0.085 0.226 0.198 0.752 0.782
@@
Pre-trained language model BERT F1 0.703 0.674 0.553 0.829 0.970
Precision 0.828 0.777 0.763 0.846 0.981
Recall 0.680 0.717 0.597 0.818 0.961
Graph models TextGCN [34] F1 0.722 0.694 0.671 0.663 0.730
@@
We also perform an ablation study with respect to the proposed phrase and phrase POS tag features. For each experiment, the feature is excluded during embedding concatenation, and the rest of the model architecture remains as it is. The results are shown in Table 7.
Table 7: Macro F1 scores obtained from ablation study
Feature Removed NU ARC LREC Bloom TREC
None (All Features) 0.724 0.712 0.751 0.672 0.801
Phrases 0.722 0.741 0.733 0.699 0.876
Phrase POS Tags 0.708 0.740 0.771 0.688 0.804
Phrases & Phrase POS Tags 0.560 0.610 0.620 0.459 0.560
6 Analysis
PQ-GCN achieves the best macro-averaged F1 scores across NU (0.724), ARC (0.712), and LREC (0.751), while consistently delivering high precision and recall. The improvement from the base SHINE model is also highlighted, as we have kept the original features as proposed in [31], showing the effectiveness of phrase-based features in improving question classification capabilities.
CNN model trained in this experiment is a relatively simple model with no specialized components, yet performs well on Bloom and TREC. This can be attributed to the characteristics of the two corpus, where presence of certain words indicate a strong association to a category. An analysis of most frequently occurring words have been included in Table 8, showing that the Bloom and TREC datasets contains distinct and mutually exclusive class-specific keywords. For example, Bloom’s Synthesis class contains high percentages of command words such as “suggest", “propose", and “design" that does not appear in questions of other classes. While the convolution kernels of CNN are effective in capturing these keywords as local features, it performs poorly on NU, ARC, and LREC where there are several overlapping non-class-specific keywords such as “would", “water", and “energy".
Table 8: Keyword Analysis for Five Datasets. Frequency is calculated by % of questions in the label set that contains the keyword.
@@ 6 Analysis
PQ-GCN performs well against BERT as well, especially for NU, ARC and LREC. However, on Bloom and TREC, BERT outperforms PQ-GCN, possibly due to the difference in the size of the dataset available for fine-tuning the model.
@@
TensorGCN performs with the highest F1 score on TREC among graph models. TensorGCN obtains its semantic embeddings by training an LSTM model on the fly during graph construction, and this method might be effective in generating more context-accurate embeddings with larger datasets. While this brings about a 0.004 point gain in F1 score over PQ-GCN, it also results in a longer graph construction time with increasing dataset size due to increased training time of the LSTM model, as compared to the proposed methodology of using pre-trained embedding models.
@@
Bi-LSTM has a consistently average performance across the datasets, which is expected of a simple sequence model. TextGCN shows apt results for NU and ARC datasets where its F1 scores are second only to PQ-GCN, but there is room for improvement in leveraging text features other than PMI and TF-IDF when constructing the graph, especially when it comes to text with niche characteristics such as questions. Text-Level GNN shows inconsistent poor performance across the datasets, as its learnable word-word edge weights are highly dependent on relative proximity of keywords and the ability of the message passing mechanism to disambiguate polysemous words based on word proximity. HyperGAT also shows inconsistent performance, possibly attributed to its word-word edge features, which are extracted by segmenting a text at sentence level and treating each sentence as a “window”, as opposed to a sliding window in PMI calculation. Questions in the datasets usually consist of one or two sentences, and lacks usual word-sentence-document hierarchy that HyperGAT leverages on. ME-GCN’s embedding distance-based edge features and InducT-GCN’s text construction that mimics that of TextGCN also do not consider question-specific features, and PQ-GCN outperforms both models in F1 scores across all datasets.
The results of the ablation study conducted in Table 7 shows that adding either phrase-based feature graph into training the model improves upon the base model, while their combined effectiveness varies across datasets. This variation is expected as the proposed approach extracts noun and verb phrases dynamically for each dataset using regex patterns, and the resulting phrase set and hence constructed graph differs for each dataset.
7 Limitations
PQ-GCN adopts a transductive approach to learning the representations, and hence it is not suited for on-line inference. While the focus of this work is to improve graph-based question classification with phrase features, we hope to inspire future work in incorporating phrase-based features into inductive methods.
We evaluate PQ-GCN only on English-based question datasets. For the proposed method to be applied to other languages, the feature extraction methods must first be adapted to syntactic rules of the target language. However, extracting regex-based noun or verb phrases may be challenging for non-configurational languages with free word order [7].
We also evaluate PQ-GCN against BERT—a representative encoder-based language model—for a more complete comparison, but not decoder-based modern large language models, as fine-tuning these models is prohibitively expensive, and generative models are not aligned with our larger focus on evaluating additive text features in embedding-based classification.
We highlight that our method is more effective in low-resource settings compared to high-resource settings. However, labeling of a large question bank is resource-intensive for educators, and they are more likely to have a smaller subset of questions to be analyzed, in which case our proposed method would be succinct.
8 Conclusion